# CGKN: A Deep Learning Framework for Modeling Complex Dynamical Systems and Efficient Data Assimilation

### v1

@@ 2.1.2 CGKN for Data Assimilation
Given the data of observed state variables $\{\mathbf{u}_{1}^{\dagger}(t_{n})\}_{n=0}^{N}$, the posterior mean of latent state variables $\{\bm{\mu}_{\mathbf{v}}(t_{n})\}_{n=0}^{N}$ in Eq. (2.3) can be estimated based on the analytic formulae in Eq. (LABEL:eq:CGKN_CGF). Similar to the latent state variables, the posterior mean $\bm{\mu}_{\mathbf{v}}$ of the latent state variables can also be transformed back to the mean estimation $\bm{\mu}$ of the original unobserved state variables $\mathbf{u}_{2}$ via the decoder $\bm{\psi}$, i.e., $\bm{\mu}(t_{n})\approx\bm{\psi}(\bm{\mu}_{\mathbf{v}}(t_{n}))$. Although this direct nonlinear transformation of the mean values is efficient, it is generally not valid for an arbitrary nonlinear decoder. However, its validity can be greatly enhanced by a trainable decoder with the DA loss being considered, and more details about the DA loss are discussed in Section 2.2. An illustration of CGKN for DA is shown in Fig. 2.2(c).
@@
Compared to ensemble-based DA methods commonly used for nonlinear models, the analytic DA formulae of the CGKN model can avoid sampling errors and reduce the computational cost. This allows the DA results of CGKN to be incorporated into the loss function during the training procedure. In each training step, the difference between the data of true unobserved state $\mathbf{u}_{2}^{\dagger}$ and the posterior mean $\bm{\mu}$ is quantified by a DA loss term, which promotes the DA performance of the trained CGKN model. More details about the DA loss term and total loss function are summarized in Section 2.2.
@@
To efficiently quantify the uncertainties associated with the posterior mean of the unobserved state $\mathbf{u}_{2}$, this work explores a post-processing tool based on residual analysis. Residual analysis [93, 94, 95] is a method used in statistical modeling and stochastic computing to assess how well a model fits the observed data. With a trained CGKN model, we can perform efficient data assimilation via Eq. (2.3) and the decoder $\bm{\psi}$ to obtain posterior mean $\{\bm{\mu}(t_{n})\}_{n=0}^{N}$ of unobserved state $\mathbf{u}_{2}$. The residual $\mathbf{r}$ is then calculated as the absolute difference between true data $\mathbf{u}_{2}^{\dagger}$ and the DA posterior mean $\bm{\mu}$, i.e., $\mathbf{r}(t_{n})=|\mathbf{u}_{2}^{\dagger}(t_{n})-\bm{\mu}(t_{n})|$, for $n=0,1,...,N$. The residual $\mathbf{r}$ is used to indicate the uncertainties associated with the posterior mean $\bm{\mu}$. In practice, the residual $\mathbf{r}$ is assumed to be a function of observed state variables $\mathbf{u}_{1}$, and the function can be approximated by an auxiliary neural network, with the standard regression task based on a dataset of $\{(\mathbf{u}^{\dagger}_{1}(t_{n}),\mathbf{r}(t_{n}))\}_{n=0}^{N}$ that are available from a trained CGKN model. Under Gaussian assumption, the output of the trained auxiliary neural network can be regarded as estimated standard deviation of the associated posterior mean from maximum likelihood estimation (MLE).
2.2 Learning CGKN Model from Data
@@
In this work, the proposed CGKN model involves several terms to be determined from data, including the encoder $\bm{\varphi}$, the decoder $\bm{\psi}$, and functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, and $\mathbf{g}_{2}$ in the conditional Gaussian nonlinear system of Eq. (2.3). As illustrated in Fig. 2.2, these unknown terms are approximated by an autoencoder that contains both the encoder and the decoder and a deep neural network for the unknown functions in the conditional Gaussian nonlinear system. To train the autoencoder and the deep neural network jointly, the total loss function consists of four types of loss terms.
@@
The first loss term is the autoencoder loss $L_{\textrm{AE}}$, which essentially ensures that the latent state variables $\mathbf{v}$ contain enough information of the original unobserved state variables $\mathbf{u}_{2}$. This is achieved by minizing the difference between the true data $\mathbf{u}_{2}^{\dagger}$ and the outputs of autoencoder $\bm{\psi}(\mathbf{v}^{\dagger})$, where $\mathbf{v}^{\dagger}=\bm{\varphi}(\mathbf{u}^{\dagger}_{2})$. It is worth noting that the tunable nonlinearity within the encoder $\bm{\varphi}$ allows the conditional linear dynamics of the latent state variables $\mathbf{v}$ in Eq. (2.3), which is inspired by the Koopman theory, even though the true dynamics of $\mathbf{u}_{2}$ can be strongly nonlinear.
@@
The second and third loss terms are the forecast loss of the original state denoted as $L_{\mathbf{u}}$ and the forecast loss of the latent state denoted as $L_{\mathbf{v}}$. The $L_{\mathbf{u}}$ quantifies the difference between true state $\mathbf{u}^{\dagger}$ and the state prediction $\mathbf{u}$, while the $L_{\mathbf{v}}$ measures the difference between true latent state $\mathbf{v}^{\dagger}=\bm{\varphi}(\mathbf{u}_{2}^{\dagger})$ and latent state prediction $\mathbf{v}$. The motivation for incorporating the forecast loss of the original state variables $\mathbf{u}$ is straightforward, as one of the applications for CGKN model is to provide state prediction of the true complex dynamical system. On the other hand, we found that the forecast loss $L_{\mathbf{v}}$ of the latent state variables $\mathbf{v}$ can enhance the predictive performance of the CGKN model, and the main reason is that the forecast loss $L_{\mathbf{v}}$ promotes the predictive capability of the conditional Gaussian nonlinear system in Eq. (2.3).
@@
The fourth loss term is the DA loss $L_{\textrm{DA}}$, which is enabled by the efficient analytic formulae in Eq. (LABEL:eq:CGKN_CGF) to obtain the posterior mean $\bm{\mu}_{\mathbf{v}}$ of the latent state variables $\mathbf{v}$. It is worth noting that the DA performance of the original unobserved state variables $\mathbf{u}_{2}$ is of actual interest. In this work, we directly map the posterior mean $\bm{\mu}_{\mathbf{v}}$ to approximate the posterior mean $\bm{\mu}$ of $\mathbf{u}_{2}$ via the decoder, i.e., $\bm{\mu}\approx\bm{\psi}(\bm{\mu}_{\mathbf{v}})$. Although such an approximation is often invalid for an arbitrary nonlinear decoder $\bm{\psi}$, we enforce the difference between the mapped posterior mean $\bm{\mu}$ and the data of $\mathbf{u}_{2}$ to be small via the DA loss term, which effectively promotes the nonlinearity within the trained decoder such that the direct mapping of posterior mean values serves as a good approximation. In addition, the DA loss can potentially enhance the discovery of probabilistic relation between the observed state variables $\mathbf{u}_{1}$ and the unobservable ones $\mathbf{u}_{2}$ when training the CGKN model.
@@
$L_{\textrm{AE}}$ $:=\mathbb{E}_{\mathbf{u}_{2}^{\dagger}\sim p(\mathbf{u}_{2}^{\dagger})}\|\mathbf{u}_{2}^{\dagger}-\bm{\psi}(\bm{\varphi}(\mathbf{u}_{2}^{\dagger}))\|^{2},$ (2.9)
$L_{\mathbf{u}}$ $:=\mathbb{E}_{\mathbf{u}^{\dagger}(t_{0})\sim p(\mathbf{u}^{\dagger}(t_{0}))}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{u}^{\dagger}(t_{i})-\mathbf{u}(t_{i})\|^{2},$ (2.10)
$L_{\mathbf{v}}$ $:=\mathbb{E}_{\mathbf{u}^{\dagger}(t_{0})\sim p(\mathbf{u}^{\dagger}(t_{0}))}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{v}^{\dagger}(t_{i})-\mathbf{v}(t_{i})\|^{2},$ (2.11)
$L_{\textrm{DA}}$ $:=\mathbb{E}_{\mathbf{u}^{\dagger}(t_{0})\sim p(\mathbf{u}^{\dagger}(t_{0}))}\frac{1}{N_{l}-N_{b}}\sum_{\begin{subarray}{c}j=N_{b}+1\end{subarray}}^{N_{l}}\|\mathbf{u}^{\dagger}_{2}(t_{j})-\bm{\mu}(t_{j})\|^{2},$ (2.12)
@@
which can be solved by standard stochastic gradient descent algorithms, and the gradient is readily available via the implementation of the CGKN model on a platform that supports automatic differentiation.
@@
$\mathrm{diag}(\bm{\sigma}_{1})=\sqrt{\frac{\Delta t}{N_{t}}\sum_{n=1}^{N_{t}}\bigl{(}\dot{\mathbf{u}}^{\dagger}_{1}(t_{n})-\dot{\mathbf{u}}_{1}(t_{n})\bigr{)}\odot\bigl{(}\dot{\mathbf{u}}_{1}^{\dagger}(t_{n})-\dot{\mathbf{u}}_{1}(t_{n})\bigr{)}},$ (2.14)
where $\Delta t$ denotes the time step of the training data, $N_{t}$ represents the total number of time steps in the training data, and the notation $\odot$ signifies element-wise multiplication. In Eq. (2.14), $\dot{\mathbf{u}}_{1}^{\dagger}(t_{n})$ is the time derivative of the true observed state variables, which can be computed by taking the numerical derivatives of $\mathbf{u}_{1}$ at time $t_{n}$, while $\dot{\mathbf{u}}_{1}(t_{n})$ is the time derivative of the corresponding prediction from the pre-trained CGKN model without the DA loss. The coefficient $\bm{\sigma}_{2}$ represents the noise associated with the latent state $\mathbf{v}$, which does not physically exist and, therefore, lacks a true value. It is also assumed to be a diagonal matrix in the form $c\mathbf{I}_{d_{\mathbf{v}}}$, where $c\in\mathbb{R}$ is a scalar and $\mathbf{I}_{d_{\mathbf{v}}}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ is the identity matrix. In addition, $\bm{\sigma}_{2}$ in the analytic formulae of Eq. (LABEL:eq:CGKN_CGF) acts as a bias term, and thus $c$ can be treated as either a trainable parameter or a constant. In this work, the scalar $c$ is manually set as a constant.

### v2

@@ 2.2.2 CGKN for Data Assimilation
Given the data of observed state variables $\{\mathbf{u}_{1}^{\star}(t_{n})\}_{n=0}^{N}$, the posterior mean of latent state variables $\{\bm{\mu}_{\mathbf{v}}(t_{n})\}_{n=0}^{N}$ in Eq. (2.8) can be estimated based on the analytic formulae in Eq. (LABEL:eq:CGKN_CGF). Similar to the latent state variables, the posterior mean $\bm{\mu}_{\mathbf{v}}$ of the latent state variables can also be transformed back to the mean estimation $\bm{\mu}$ of the original unobserved state variables $\mathbf{u}_{2}$ via the decoder $\bm{\psi}$, i.e., $\bm{\mu}(t_{n})\approx\bm{\psi}(\bm{\mu}_{\mathbf{v}}(t_{n}))$. Although this direct nonlinear transformation of the mean values is efficient, it is generally not valid for an arbitrary nonlinear decoder. However, its validity can be greatly enhanced by a trainable decoder with the DA loss being considered. In the DA process, the exact initial condition or initial distribution of the DA solution is usually unknown. As a result, it takes some time for the DA solution to stabilize and adjust to eliminate the inconsistency from the initialization. During this initial period of DA, which is known as the warm-up period, the estimated states may deviate from the actual values. Consequently, warm-up period is crucial for ensuring that DA produces stable and accurate estimates for unobserved states before proceeding with any analysis or decision-making. Therefore, the steps of DA for the warm-up period are omitted in the evaluation of DA loss for training the CGKN. More details about the DA loss are discussed in Section 2.3. An illustration of CGKN for DA is shown in Fig. 2.2(c).
The effectiveness of DA depends on the observability of unobserved states, which is the intrinsic property of the complex system described in Eq. (LABEL:eq:DS_u1u2). The observability of unobserved states is determined by the dynamics functions $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. For instance, $\mathbf{u}_{2}$ is fully observable if the observed states $\mathbf{u}_{1}$ sufficiently dictate the dynamics of $\mathbf{u}_{2}$ in such a way that observing $\mathbf{u}_{1}$ over time gives all the necessary information to estimate $\mathbf{u}_{2}$ without ambiguity. If the $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are linear functions, the observability matrix condition can be used to determine the full observability. $\mathbf{u}_{2}$ is not observable if the dynamics function $\mathcal{M}_{1}$ of the observed state does not include $\mathbf{u}_{2}$. In this case, the observation of $\mathbf{u}_{1}$ does not provide any information to infer the unobserved state $\mathbf{u}_{2}$, and consequently, CGKN will be unable to accurately predict future states using the estimated unobserved states from DA.
@@
Efficient DA can be performed by CGKN with the analytic formulae in Eq. (LABEL:eq:CGKN_CGF). The computational complexity for this analytic approach is $\mathcal{O}(N^{3})$ where $N=d_{\mathbf{v}}$ is the dimension of the latent state. As a comparison, the computational complexity of ensemble DA methods is $\mathcal{O}(JM^{2})$, where $J$ is the ensemble size and $M=d_{\mathbf{u}_{2}}$ is the dimension of unobserved states. To accurately approximate the probabilistic distribution of an $N$-dimensional random vector, an ensemble size $J=c^{M}$ is typically required, where $c$ denotes the sample size for a single random variable. This leads to a resulting complexity of $\mathcal{O}(c^{M}M^{2})$, which grows exponentially due to the curse of dimensionality. In real-world applications, a much smaller ensemble size is often adopted due to the restriction of computational resources, which inevitably leads to issues originating from sampling errors, and various techniques (e.g., localization) have been developed to empirically deal with insufficient sample size in a high-dimensional space. Therefore, compared with ensemble-based DA methods, which are commonly used for nonlinear models, the analytic DA formulae of the CGKN model can avoid sampling errors and reduce the computational cost. This allows the DA results of CGKN to be incorporated into the loss function during the training procedure. In each training step, the difference between the data of true unobserved state $\mathbf{u}_{2}^{\star}$ and the posterior mean $\bm{\mu}$ is quantified by a DA loss term, which promotes the DA performance of the trained CGKN model. More details about the DA loss term and total loss function are summarized in Section 2.3.
@@
To efficiently quantify the uncertainties associated with the posterior mean of the unobserved state $\mathbf{u}_{2}$, this work explores a post-processing tool based on residual analysis. Residual analysis [96, 97, 98] is a method used in statistical modeling and stochastic computing to assess how well a model fits the observed data. With a trained CGKN model, we can perform efficient data assimilation via Eq. (2.8) and the decoder $\bm{\psi}$ to obtain posterior mean $\{\bm{\mu}(t_{n})\}_{n=0}^{N}$ of unobserved state $\mathbf{u}_{2}$. The residual $\mathbf{r}$ is then calculated as the absolute difference between true data $\mathbf{u}_{2}^{\star}$ and the DA posterior mean $\bm{\mu}$, i.e., $\mathbf{r}(t_{n})=|\mathbf{u}_{2}^{\star}(t_{n})-\bm{\mu}(t_{n})|$, for $n=0,1,...,N$. The residual $\mathbf{r}$ is used to indicate the uncertainties associated with the posterior mean $\bm{\mu}$. In practice, the residual $\mathbf{r}$ is assumed to be a function of observed state variables $\mathbf{u}_{1}$, and the function can be approximated by an auxiliary neural network, with the standard regression task based on a dataset of $\{(\mathbf{u}^{\star}_{1}(t_{n}),\mathbf{r}(t_{n}))\}_{n=0}^{N}$ that are available from a trained CGKN model. Under Gaussian assumption, the output of the trained auxiliary neural network can be regarded as estimated standard deviation of the associated posterior mean from maximum likelihood estimation (MLE).
2.3 Learning CGKN from Data
@@
In this work, the proposed CGKN model involves several terms to be determined from data, including the encoder $\bm{\varphi}$, the decoder $\bm{\psi}$, and functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, and $\mathbf{g}_{2}$ in the conditional Gaussian nonlinear system of Eq. (2.8). As illustrated in Fig. 2.2, these unknown terms are approximated by an autoencoder that contains both the encoder and the decoder and deep neural networks for the unknown functions in the conditional Gaussian nonlinear system. To train the autoencoder and the deep neural networks jointly, the total loss function consists of four types of loss terms.
@@
The first loss term is the autoencoder loss $L_{\textrm{AE}}$, which essentially ensures that the latent state variables $\mathbf{v}$ contain enough information of the original unobserved state variables $\mathbf{u}_{2}$. This is achieved by minimizing the difference between the true data $\mathbf{u}_{2}^{\star}$ and the outputs of autoencoder $\bm{\psi}(\mathbf{v}^{\star})$, where $\mathbf{v}^{\star}=\bm{\varphi}(\mathbf{u}^{\star}_{2})$. It is worth noting that the tunable nonlinearity within the encoder $\bm{\varphi}$ allows the conditional linear dynamics of the latent state variables $\mathbf{v}$ in Eq. (2.8), which is inspired by the Koopman theory, even though the true dynamics of $\mathbf{u}_{2}$ can be strongly nonlinear. The autoencoder loss $L_{\textrm{AE}}$ is defined as $\mathbb{E}_{\mathbf{u}_{2}^{\star}}\|\mathbf{u}_{2}^{\star}-\bm{\psi}(\bm{\varphi}(\mathbf{u}_{2}^{\star}))\|^{2}.$
@@
The second and third loss terms are the forecast loss of the original state denoted as $L_{\mathbf{u}}$ and the forecast loss of the latent state denoted as $L_{\mathbf{v}}$. The $L_{\mathbf{u}}$ quantifies the difference between true state $\mathbf{u}^{\star}$ and the state prediction $\mathbf{u}$, while the $L_{\mathbf{v}}$ measures the difference between true latent state $\mathbf{v}^{\star}=\bm{\varphi}(\mathbf{u}_{2}^{\star})$ and latent state prediction $\mathbf{v}$. The motivation for incorporating the forecast loss of the original state variables $\mathbf{u}$ is straightforward, as one of the applications for CGKN model is to provide state prediction of the true complex dynamical system. On the other hand, we found that the forecast loss $L_{\mathbf{v}}$ of the latent state variables $\mathbf{v}$ can enhance the predictive performance of the CGKN model, and the main reason is that the forecast loss $L_{\mathbf{v}}$ promotes the predictive capability of the conditional Gaussian nonlinear system in Eq. (2.8). The formulas of forecast loss of the original state $L_{\mathbf{u}}$ and the forecast loss of the latent state $L_{\mathbf{v}}$ are $\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{u}^{\star}(t_{i})-\mathbf{u}(t_{i})\|^{2}$ and $\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{v}^{\star}(t_{i})-\mathbf{v}(t_{i})\|^{2}$, respectively, where $N_{s}$ represents the number of forecast steps.
@@
The fourth loss term is the DA loss $L_{\textrm{DA}}$, which is enabled by the efficient analytic formulae in Eq. (LABEL:eq:CGKN_CGF) to obtain the posterior mean $\bm{\mu}_{\mathbf{v}}$ of the latent state variables $\mathbf{v}$. It is worth noting that the DA performance of the original unobserved state variables $\mathbf{u}_{2}$ is of actual interest. In this work, we directly map the posterior mean $\bm{\mu}_{\mathbf{v}}$ to approximate the posterior mean $\bm{\mu}$ of $\mathbf{u}_{2}$ via the decoder, i.e., $\bm{\mu}\approx\bm{\psi}(\bm{\mu}_{\mathbf{v}})$. Although such an approximation is often invalid for an arbitrary nonlinear decoder $\bm{\psi}$, we enforce the difference between the mapped posterior mean $\bm{\mu}$ and the data of $\mathbf{u}_{2}$ to be small via the DA loss term, which effectively promotes the nonlinearity within the trained decoder such that the direct mapping of posterior mean values serves as a good approximation. In addition, the DA loss can potentially enhance the discovery of probabilistic relation between the observed state variables $\mathbf{u}_{1}$ and the unobserved ones $\mathbf{u}_{2}$ when training the CGKN model. The DA loss $L_{\textrm{DA}}$ is defined as $\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{l}-N_{b}}\sum_{\begin{subarray}{c}j=N_{b}+1\end{subarray}}^{N_{l}}\|\mathbf{u}^{\star}_{2}(t_{j})-\bm{\mu}(t_{j})\|^{2}$, where $N_{l}$ is the number of DA steps and $N_{b}$ is the warm-up period steps.
@@
$L_{\textrm{AE}}$ $:=\mathbb{E}_{\mathbf{u}_{2}^{\star}}\|\mathbf{u}_{2}^{\star}-\bm{\psi}(\bm{\varphi}(\mathbf{u}_{2}^{\star}))\|^{2},$ (2.14)
$L_{\mathbf{u}}$ $:=\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{u}^{\star}(t_{i})-\mathbf{u}(t_{i})\|^{2},$ (2.15)
$L_{\mathbf{v}}$ $:=\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\|\mathbf{v}^{\star}(t_{i})-\mathbf{v}(t_{i})\|^{2},$ (2.16)
$L_{\textrm{DA}}$ $:=\mathbb{E}_{\mathbf{u}^{\star}(t_{0})}\frac{1}{N_{l}-N_{b}}\sum_{\begin{subarray}{c}j=N_{b}+1\end{subarray}}^{N_{l}}\|\mathbf{u}^{\star}_{2}(t_{j})-\bm{\mu}(t_{j})\|^{2},$ (2.17)
@@
which can be solved by standard stochastic gradient descent algorithms, and the gradient is readily available via the implementation of the CGKN model on a platform that supports automatic differentiation. A common choice for the gains is $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=1/d_{\mathbf{u}}$, $\lambda_{\mathbf{v}}=1/d_{\mathbf{u}}$, and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$, which scale each loss function by a factor of the inverse of the respective vector’s dimension. The resulting loss functions are the mean squared error (MSE) between the true data and the approximated data. In practice, the gains can also be set as hyper-parameters or manually adjusted based on the domain knowledge of users.
@@
$\mathrm{diag}(\bm{\sigma}_{1})=\sqrt{\frac{\Delta t}{N_{t}}\sum_{n=1}^{N_{t}}\bigl{(}\dot{\mathbf{u}}^{\star}_{1}(t_{n})-\dot{\mathbf{u}}_{1}(t_{n})\bigr{)}\odot\bigl{(}\dot{\mathbf{u}}_{1}^{\star}(t_{n})-\dot{\mathbf{u}}_{1}(t_{n})\bigr{)}},$ (2.19)
where $\Delta t$ denotes the time step of the training data, $N_{t}$ represents the total number of time steps in the training data, and the notation $\odot$ signifies element-wise multiplication. In Eq. (2.19), $\dot{\mathbf{u}}_{1}^{\star}(t_{n})$ is the time derivative of the true observed state variables, which can be computed by taking the numerical derivatives of $\mathbf{u}_{1}$ at time $t_{n}$, while $\dot{\mathbf{u}}_{1}(t_{n})$ is the time derivative of the corresponding prediction from the pre-trained CGKN model without the DA loss. The coefficient $\bm{\sigma}_{2}$ represents the noise associated with the latent state $\mathbf{v}$, which does not physically exist and, therefore, lacks a true value. It is also assumed to be a diagonal matrix in the form $c\mathbf{I}_{d_{\mathbf{v}}}$, where $c\in\mathbb{R}$ is a scalar and $\mathbf{I}_{d_{\mathbf{v}}}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ is the identity matrix. In addition, $\bm{\sigma}_{2}$ in the analytic formulae of Eq. (LABEL:eq:CGKN_CGF) acts as a bias term, and thus $c$ can be treated as either a trainable parameter or a constant. In this work, the scalar $c$ is manually set as a constant.